1..calculate the simple interest on a loan of $18.500 at 11% annual interest for 18months.
2..an amusement park purchased equipment from a manufacturer for a list price 42,000,minus trade discounts of 30/20/15. calculate the amount of trade discount on the purchase.
3...a retail jewler sells a pair of gold earrings for 213.90. the jewler uses a 130% markup based on cost. what was thw wholesale cost of the earrings.
4..what would be the amount of compound interest on $6,000 invested for two years at 9% compounded semiannually?

Answer by rfer(16322)

1)
I=Prt
I=18500(.11)(1.5)=3052.50
2)
42000(.7)=29400
29400(.8)=23520
23520(.85)=19992
42000-19992=$22,008 discount
3)
213.90/2.30=$93
4)
A=P(1+r/n)^nt
A=6000(1+.09/2)^2*2
A=6000(1.045)^4
A=6000(1.1925186)
A=7155.11
7155.11-6000=$1155.11 interest